# Conceptualization, Methodology, Investigation, Data curation, Formal analysis, Writing - Original draft preparation

### v4

@@ Abstract
Security code review is a time-consuming and labor-intensive process typically requiring integration with automated security defect detection tools. However, existing security analysis tools struggle with poor generalization, high false positive rates, and coarse detection granularity. Large Language Models (LLMs) have been considered promising candidates for addressing those challenges. In this study, we conducted an empirical study to explore the potential of LLMs in detecting security defects during code review. Specifically, we evaluated the performance of six LLMs under five different prompts and compared them with state-of-the-art static analysis tools. We also performed linguistic and regression analyses for the best-performing LLM to identify quality problems in its responses and factors influencing its performance. Our findings show that: (1) existing pre-trained LLMs have limited capability in security code review but significantly outperform the state-of-the-art static analysis tools. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) GPT-4 frequently generates verbose or non-compliant responses with the task requirements given in the prompts. (4) GPT-4 is more adept at identifying security defects in code files with fewer tokens, containing functional logic, or written by developers with less involvement in the project.
@@ 1 Introduction
Security defects are potential security risks or weaknesses in the system introduced during implementation (Paul et al., 2021). Once attackers maliciously exploit a security defect, it could lead to serious consequences, including data breaches, financial losses, and service disruptions (Telang and Wattal, 2007; Cavusoglu et al., 2004). The longer a security defect stays in the program, the higher its associated fixing and maintenance costs are (Planning, 2002). Thus, many organizations are shifting security analysis to earlier stages of software development, typically at code review time (GitLab, 2022). Code review is a human-intensive process in which the reviewer examines the code submitted by the developer to detect bugs, verify the implementation of the specification, ensure compliance with guidelines, and ensure quality. During this process, the reviewer raises issues in the code, discusses them with the developer, and provides recommendations. Adopting security analysis in code review can incorporate diverse viewpoints from both reviewers and developers, and effectively prevent security defects created by programmers working alone, as they naturally have a limited perspective of potential security risks (Weir et al., 2017), from being merged into the source code repository. In this study, we focus on security analysis in the context of code review, referring to it as security code review (Edmundson et al., 2013).
@@
Security code review is increasingly integrated into the development pipelines by project teams (Paul et al., 2021). Its resource-intensive nature, which requires significant human effort and time to review and revise the code, poses a notable challenge, particularly in popular large-scale open-source projects with numerous contributions (Guo et al., 2024). Therefore, it has been desired to develop effective automated tools to assist code reviewers in security code review (Gholami and Amri, 2021). Despite various static/dynamic/hybrid program analysis tools being proposed, they each encounter practical challenges, such as excessive imprecision (Nong et al., 2021), high false positive rates (Johnson et al., 2013; Sui et al., 2020), input range limitations (Haller et al., 2013), and weak scalability (Nong et al., 2021).
@@
Several data-driven approaches based on machine learning (ML)/deep learning (DL) have been developed for vulnerability detection and repair, which can somewhat handle the challenges encountered by traditional program analysis tools (Chakraborty et al., 2021). However, due to the lack of sizable and realistic training datasets (Nong et al., 2022), those approaches inevitably lacked robustness and tended to fall short when implemented in unfamiliar, real-world projects (Nong et al., 2024a). Nong et al. (2024b) found that even using the training dataset augmented by the latest vulnerability injection technique, the highest F1 score and the top-1 accuracy that state-of-the-art DL-based vulnerability detection and repair tools can achieve on the dataset constructed from real-world projects are still limited, at 20.01% and 21.05% respectively. Therefore, it remains essential to further explore techniques with better generalization and robustness to address vulnerabilities in real-world projects.
@@
Recently, pre-trained LLMs have demonstrated promising performances across a broad spectrum of software engineering tasks, such as program repair (Jiang et al., 2023), test generation (Kang et al., 2023; Schäfer et al., 2023), and specification generation (Xie et al., 2023). Such models can understand and generate explanatory natural language responses, and are thus considered a potential candidate for security code review. We expect that LLM can act as a code reviewer, not only identifying security defects in code files, but also explaining the specific defect scenarios in detail to developers in a human-readable way. Applying LLMs to support security code review can help detect security defects earlier and improve productivity by helping code reviewers check their code faster and more efficiently. However, applying LLMs in the specific field of security code review remains largely unexplored.
@@
Many studies have focused on leveraging LLMs for vulnerability detection. Zhou et al. (2024b) evaluated the capabilities of GPT-3.5 and GPT-4 in vulnerability detection under few-shot learning prompts, while Purba et al. (2023) compared LLMs with static analysis tools in detecting software vulnerabilities. These studies mainly conducted a coarse-grained assessment of LLMs in binary classification tasks to judge whether there were security defects in the code. They did not require LLMs to provide more detailed information, such as the location and type of vulnerability. Researchers have also developed prompting strategies to enhance the effectiveness of LLMs in vulnerability detection, such as the DL-based prompting framework proposed by Yang et al. (2024) and the vulnerability-semantic guided prompting approach formulated by Nong et al. (2024a). However, the datasets utilized in these studies are synthetic and CVE datasets, which do not fully represent real-world codebases.
@@
Motivated by the limitations mentioned above, we aim to bridge the knowledge gap by conducting an empirical study to comprehensively explore the potential of LLMs in fine-grained security code review on a practical dataset constructed from real-world code repositories. Specifically, in this study, LLMs were requested to provide detailed information on the identified security defect, including its line number, type, specific description, and suggested fix. The dataset we utilized consists of 534 code review files obtained from four open-source projects (namely, OpenStack Nova and Neutron, and Qt Base and Creator). We evaluated the performance of LLMs, analyzed existing quality problems, and examined factors influencing their performance to offer insights into the real-world applicability and limitations of LLM in security code review. First, we identified five prompt templates to enhance LLMs (based on the prompting strategies formulated by Zhang et al. (2023)). Then, we compared the performance of 6 open-source and closed-source LLMs with static analysis tools across these prompt templates in detecting security defects identified by reviewers. After selecting the responses generated by the best-performing LLM-prompt combination, we first manually extracted quality problems from these responses. We then used these responses to construct a cumulative link mixed model to explore the impact of 10 factors on the performance of the LLM. Our findings reveal that: (1) current popular general-purpose LLMs exhibit limited capability in detecting security defects identified by reviewers in code reviews. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) Inconcise and non-compliant responses to prompts are common problems encountered by GPT-4 in security code review. (4) GPT-4 is more likely to identify security defects in code files with fewer tokens, containing functional logic, and written by developers with less involvement in the project. In summary, this study makes the following contributions:
(1) We conducted a fine-grained evaluation of the capability of the currently popular LLMs in security code review. (2) We measured the impact of the non-determinism of LLMs on the consistency of security code review. (3) We identified quality problems and quantified hallucinations in the responses generated by the top-performing LLM, aiming to highlight the challenges faced by LLMs in security code review. (4) We carried out the first analysis of the factors that may influence the performance of the LLM in security code review.
@@
Paper Organization. Section 6 surveys the related work. Section 2 describes the methodology employed in this study. Section 3 presents the results, followed by a discussion of the implications of our study results in Section 4. Section 5 clarifies the threats to validity. Finally, Section 7 concludes this work and outlines future directions. The replication package of this study has been provided online (Yu et al., 2025).
@@ 2.1 Research Questions
Motivation. Many studies have focused on the performance of LLMs in detecting security defects, but they often suffer from low-quality datasets collected automatically and a narrow detection scope primarily confined to function-level code (Zhou et al., 2024a). To address these limitations, RQ1 utilizes a manually curated dataset to ensure data quality. It comprehensively evaluates the capability of LLMs to review a complete code file from real-world code reviews for security defects. Additionally, due to the inherent non-determinism of LLMs, the responses of LLMs may exhibit considerable divergence for the same prompt across different runs. Non-determinism is mainly caused by decoding configurations (Song et al., 2024), with other hypotheses such as floating point, unreliable GPU calculations (Ouyang et al., 2025), and its sparse MoE architecture failing to enforce per-sequence determinism (La Malfa et al., 2024; Puigcerver et al., 2023). Considering that the non-determinism of LLMs has not yet been studied when applied to security tasks, we conducted a quantitative analysis of the consistency of LLM responses to explore its impact on the task of security code review.
@@
RQ2.2: Which of these quality problems are hallucination-related, and what are their proportions?
Motivation. Kabir et al. (2024) noted that LLMs’ responses may be verbose, inconsistent, and contain conceptual or logical errors, negatively impacting the quality of the answers. Among these problems, hallucination in particular has been extensively studied  (Liu et al., 2024; Tonmoy et al., 2024). Hallucination refers to content generated by LLMs that appears plausible but lacks factual grounding (Huang et al., 2025). If we applied LLMs to real-world software development tasks, particularly during security code review, such hallucination may mislead practitioners, increase review costs, and even compromise system security. Previous studies have not yet thoroughly examined quality problems present in LLM-generated responses during security code review. RQ2 intends to conduct an in-depth linguistic analysis to investigate possible hallucination issues. First, we investigated the distribution of quality problems in LLMs’ responses, such as Vague statement, Incorrect concept, and Incomplete answer. Then we identified hallucination-related issues among these problems and quantified their proportion to bridge the knowledge gap and highlight potential areas for improvement.
@@
Motivation. Previous studies have identified potential factors influencing the effectiveness of LLMs in detecting security defects (Chen et al., 2023; Kabir et al., 2024; Zhou et al., 2023), primarily through case studies or observations. However, there is a lack of evidence substantiating such claims. RQ3 employs a regression analysis to validate whether these factors significantly impact the performance of LLMs. This RQ helps pinpoint areas of concern in applying LLMs to security code review and highlights possible improvement directions.
@@ 2.2 Data Collection
We leveraged the code review dataset constructed by Yu et al. (2023) and examined the capability of LLMs in security code review. The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types from four projects: OpenStack Nova and Neutron (almost entirely written in Python), and Qt Base and Creator (mostly in C++ with some portions in C, Objective-C++, and Python). These projects are the most active within OpenStack and Qt (Hirao et al., 2020) and have been widely used in code review-related studies (Han et al., 2022; Hamasaki et al., 2013; Spadini et al., 2018). For each security defect, we could not include the code from the entire patchset in our prompt, which is always too long and exceeds the input token limits of LLMs. Instead, we included only the code from the single file, where the reviewer commented to identify a security defect. The 614 security-related review comments were reorganized as follows:
@@
Our final dataset consists of 534 code files (258 Python and 276 C/C++ files), each including the source code, the commit message of the associated patchset, and detailed information of the target security defects as identified by the reviewers.
@@ 2.3 Research Procedure of RQ1
Model Version Ctx Len Parameter
temperature top_p
@@
GPT-4 Turbo gpt-4-1106-preview 128k 1.0 1.0
Gemini Pro gemini-1.0-pro 32k 0.9 1.0
Llama 2 7B llama-2-7b-chat 4k 1.0 0.9
Llama 2 70B llama-2-70b-chat 4k 1.0 0.9
@@ 2.3.1 LLM Selection
To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December 2023. Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through the API. For open-source LLMs, we deployed the smallest and largest chat versions of models in the Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs, each with 24 GB of VRAM. To realistically reflect on the practical capabilities of LLMs in security defect detection, we used the default hyperparameter settings throughout our experiments. Some studies mitigate the non-determinism of LLMs by setting the temperature to zero (Atil et al., 2024; Blackwell et al., 2024). In contrast, we refrained from doing so, as this may adversely affect the performance of the LLMs (Song et al., 2024) and may not guarantee determinism (Ouyang et al., 2025; Astekin et al., 2024). We retained the default temperature settings and quantified the non-determinism of LLMs as one aspect of performance evaluation in RQ1.4, rather than artificially suppressing such non-determinism via hyperparameter tuning. The versions, context lengths (Ctx Len), and parameter settings of LLMs used in our study are summarized in Table 1.
Notably, GPT-3.5 and GPT-4 can be accessed through both API and the ChatGPT platform. During our preliminary exploration stage, we compared the performances of GPT-4 and GPT-3.5 accessed by these two ways. The two models provided by the ChatGPT platform performed better and thus were employed in our study.

### v5

@@ Abstract
Security code review is a time-consuming and labor-intensive process typically requiring integration with automated security defect detection tools. However, existing security analysis tools struggle with poor generalization, high false positive rates, and coarse detection granularity. Large Language Models (LLMs) have been considered promising candidates for addressing those challenges. In this study, we conducted an empirical study to explore the potential of LLMs in detecting security defects during code review. Specifically, we evaluated the performance of seven LLMs under five different prompts and compared them with state-of-the-art static analysis tools. We also performed linguistic and regression analyses for the two top-performing LLMs to identify quality problems in their responses and factors influencing their performance. Our findings show that: (1) In security code review, LLMs significantly outperform state-of-the-art static analysis tools, and the reasoning-optimized LLM performs better than general-purpose LLMs. (2) DeepSeek-R1 achieves the highest performance, followed by GPT-4. The optimal prompt for DeepSeek-R1 incorporates both the commit message and chain-of-thought (CoT) guidance, while for GPT-4, the prompt with a Common Weakness Enumeration (CWE) list works best. (3) GPT-4 frequently produces vague expressions and exhibits difficulties in accurately following instructions in the prompts, while DeepSeek-R1 more commonly generates inaccurate code details in its outputs. (4) LLMs are more adept at identifying security defects in code files that have fewer tokens and security-relevant annotations.
@@ 1 Introduction
Security defects are potential security risks or weaknesses in the system introduced during implementation (paul2021security). Once attackers maliciously exploit a security defect, it could lead to serious consequences, including data breaches, financial losses, and service disruptions (telang2007empirical; cavusoglu2004effect). The longer a security defect stays in the program, the higher its associated fixing and maintenance costs are (planning2002economic). Thus, many organizations are shifting security analysis to earlier stages of software development, typically at code review time (gitlab2022url). Code review is a human-intensive process in which the reviewer examines the code submitted by the developer to detect bugs, verify the implementation of the specification, ensure compliance with guidelines, and ensure quality. During this process, the reviewer raises issues in the code, discusses them with the developer, and provides recommendations. Adopting security analysis in code review can incorporate diverse viewpoints from both reviewers and developers, and effectively prevent security defects created by programmers working alone, as they naturally have a limited perspective of potential security risks (weir2017d), from being merged into the source code repository. In this study, we focus on security analysis in the context of code review, referring to it as security code review (edmundson2013empirical).
@@
Security code review is increasingly integrated into the development pipelines by project teams (paul2021security). Its resource-intensive nature, which requires significant human effort and time to review and revise the code, poses a notable challenge, particularly in popular large-scale open-source projects with numerous contributions (guo2024exploring). Therefore, it has been desired to develop effective automated tools to assist code reviewers in security code review. Despite various static/dynamic/hybrid program analysis tools being proposed, they each encounter practical challenges, such as excessive imprecision (nong2021evaluating), high false positive rates (johnson2013don; sui2020recall), input range limitations (haller2013dowsing), and weak scalability (nong2021evaluating).
@@
Several data-driven approaches based on machine learning (ML)/deep learning (DL) have been developed for vulnerability detection and repair, which can somewhat handle the challenges encountered by traditional program analysis tools (chakraborty2021deep). However, due to the lack of sizable and realistic training datasets (nong2022generating), those approaches inevitably lacked robustness and tended to fall short when implemented in unfamiliar, real-world projects (nong2024chain). nong2024vgx found that even using the training dataset augmented by the latest vulnerability injection technique, the highest F1 score and the top-1 accuracy that state-of-the-art DL-based vulnerability detection and repair tools can achieve on the dataset constructed from real-world projects are still limited, at 20.01% and 21.05% respectively. Therefore, it remains essential to further explore techniques with better generalization and robustness to address vulnerabilities in real-world projects.
@@
Recently, pre-trained LLMs have demonstrated promising performances across a broad spectrum of software engineering tasks, such as program repair (jiang2023impact), test generation (kang2023large; schafer2023empirical), and specification generation (xie2023impact). Such models can understand and generate explanatory natural language responses, and are thus considered a potential candidate for security code review. We expect that LLM can act as a code reviewer, not only identifying security defects in code files, but also explaining the specific defect scenarios in detail to developers in a human-readable way. Applying LLMs to support security code review can help detect security defects earlier and improve productivity by helping code reviewers check their code faster and more efficiently. However, applying LLMs in the specific field of security code review remains largely unexplored.
@@
Many studies have focused on leveraging LLMs for vulnerability detection. zhou2024largelanguage evaluated the capabilities of GPT-3.5 and GPT-4 in vulnerability detection under few-shot learning prompts, while purba2023software compared LLMs with static analysis tools in detecting software vulnerabilities. These studies mainly conducted a coarse-grained assessment of LLMs in binary classification tasks to judge whether there were security defects in the code. They did not require LLMs to provide more detailed information, such as the location and type of vulnerability. Researchers have also developed prompting strategies to enhance the effectiveness of LLMs in vulnerability detection, such as the DL-based prompting framework proposed by yang2024dlap and the vulnerability-semantic guided prompting approach formulated by nong2024chain. However, the datasets utilized in these studies are synthetic and CVE datasets, which do not fully represent real-world codebases.
@@
Motivated by the limitations mentioned above, we aim to bridge the knowledge gap by conducting an empirical study to comprehensively explore the potential of LLMs in fine-grained security code review on a practical dataset constructed from real-world code repositories. Specifically, in this study, LLMs were requested to provide detailed information on the identified security defect, including its line number, type, specific description, and suggested fix. The dataset we utilized consists of 534 code review files obtained from four open-source projects (namely, OpenStack Nova and Neutron, and Qt Base and Creator). We evaluated the performance of LLMs, analyzed existing quality problems, and examined factors influencing their performance to offer insights into the real-world applicability and limitations of LLM in security code review. First, we identified five prompt templates to enhance LLMs (based on the prompting strategies formulated by zhang2024prompt). Then, we compared the performance of 7 open-source and closed-source LLMs with static analysis tools across these prompt templates in detecting security defects identified by reviewers. After collecting the responses generated by the two best-performing LLMs with their respective optimal prompts, we manually extracted quality problems present in these responses. We also used these responses to construct two separate cumulative link mixed models to explore the impact of 11 factors on the performance of each LLM. Our findings reveal that: (1) LLMs outperform state-of-the-art static analysis tools in security code review. Among LLMs, the reasoning-optimized model surpasses general-purpose models. (2) DeepSeek-R1 is the best-performing LLM, followed by GPT-4. For DeepSeek-R1, the prompt including the commit message and CoT instruction is the most effective, while for GPT-4, the prompt with a CWE list works best. (3) GPT-4 struggles with vague expressions and poor instruction-following, while DeepSeek-R1 is more prone to producing incorrect code details in its responses. (4) LLMs detect security defects more effectively in code files with fewer tokens and security-related annotations. In summary, this study makes the following contributions: (1) We conducted a fine-grained evaluation of the capability of popular and representative LLMs in security code review. (2) We measured the impact of the non-determinism of LLMs on the consistency of security code review. (3) We identified quality problems and quantified hallucinations in the responses generated by the two top-performing LLMs, aiming to highlight the challenges faced by LLMs in security code review. (4) We carried out the first analysis of the factors that might influence the performance of LLMs in security code review.
@@
Paper Organization. Section 6 surveys the related work. Section 2 describes the methodology employed in this study. Section 3 presents the results, followed by a discussion of the implications of our study results in Section 4. Section 5 clarifies the threats to validity. Finally, Section 7 concludes this work and outlines future directions. The replication package of this study has been provided online (replpack).
@@ 2.1 Research Questions
Motivation. Many studies have focused on the performance of LLMs in detecting security defects, but they often suffer from low-quality datasets collected automatically and a narrow detection scope primarily confined to function-level code (zhou2025large). To address these limitations, RQ1 utilizes a manually curated dataset to ensure data quality. It comprehensively evaluates the capability of LLMs to review a complete code file from real-world code reviews for security defects. Additionally, due to the inherent non-determinism of LLMs, the responses of LLMs may exhibit considerable divergence for the same prompt across different runs. Non-determinism is mainly caused by decoding configurations (song2024good), with other hypotheses such as floating point, unreliable GPU calculations (ouyang2025empirical), and its sparse MoE architecture failing to enforce per-sequence determinism (la2024language; puigcerver2023sparse). Considering that the non-determinism of LLMs has not yet been studied when applied to security tasks, we conducted a quantitative analysis of the consistency of LLM responses to explore its impact on the task of security code review.
@@
RQ2.2: Which of these problems are hallucination-related, and what are their proportions?
Motivation. kabir2023answers noted that LLMs’ responses may be verbose, inconsistent, and contain conceptual or logical errors, negatively impacting the quality of the answers. Among these problems, hallucination in particular has been extensively studied  (liu2024exploring; tonmoy2024comprehensive). Hallucination refers to content generated by LLMs that appears plausible but lacks factual grounding (huang2025survey). If we applied LLMs to real-world software development tasks, particularly during security code review, such hallucination may mislead practitioners, increase review costs, and even compromise system security. Previous studies have not yet thoroughly examined quality problems present in LLM-generated responses during security code review. RQ2 intends to conduct an in-depth linguistic analysis to investigate possible hallucination issues. First, we investigated the distribution of quality problems in LLMs’ responses, such as Vague statement, Incorrect concept, and Incomplete answer. Then we identified hallucination-related issues among these problems and quantified their proportion to bridge the knowledge gap and highlight potential areas for improvement.
@@
Motivation. Previous studies have identified potential factors influencing the effectiveness of LLMs in detecting security defects (chen2023chatgpt; kabir2023answers; zhout2023devil), primarily through case studies or observations. However, there is a lack of evidence substantiating such claims. RQ3 employs a regression analysis to validate whether these factors significantly impact the performance of LLMs. This RQ helps pinpoint areas of concern in applying LLMs to security code review and highlights possible improvement directions.
Figure 1: Distribution of LoC of the code files with security defects
@@
Figure 3: Distribution of security defect types on the Python and C/C++ dataset
@@ 2.2 Data Collection
We leveraged the code review dataset constructed by yu2023security and examined the capability of LLMs in security code review. The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types from four projects: OpenStack Nova and Neutron (almost entirely written in Python), and Qt Base and Creator (mostly in C++ with some portions in C, Objective-C++, and Python). These projects are the most active within OpenStack and Qt (hirao2020code) and have been widely used in code review-related studies (han2022code; hamasaki2013does; spadini2018testing). For each security defect, we could not include the code from the entire patchset in our prompt, which is always too long and exceeds the input token limits of LLMs. Instead, we included only the code from the single file, where the reviewer commented to identify a security defect. The 614 security-related review comments were reorganized as follows:
@@
Our final dataset consists of 534 code files (258 Python and 276 C/C++ files), each including the source code, the commit message of the associated patchset, and detailed information of the target security defects as identified by the reviewers. We analyzed the distribution of security defect types in Python and C/C++ code files (see Fig. 3). In Python code files, the Python language employs automatic memory management and garbage collection. Therefore, Memory Leak and Buffer Overflow were relatively rare, while Race Condition was the most common type, accounting for 65.1%. In contrast, C/C++ requires programmers to manage memory and pointers manually, thus the C/C++ dataset has a more balanced distribution of security defects. In C/C++ code files, Crash was the most prevalent (32%), followed by Race Condition (17.4%), Resource Leak (16.3%) and Integer Overflow (14.9%). We also counted the lines of code (LoC) of the collected code files in our dataset. Fig. 1 shows the distribution of LoC for different languages (Python and C/C++, respectively) and the distribution of LoC for the entire dataset. We found that the median LoC is 815 and the average reaches 2,107, which substantially exceeds the scale of function-level datasets like Big-Vul (avg. 30 LoC) and SVEN (avg. 168 LoC) adopted in related studies (yin2024multitask; steenhoek2024err).
@@ 2.3 Research Procedure of RQ1
Model Version Context Window Parameter
temperature top_p Other parameters
@@
GPT-4 Turbo gpt-4-1106-preview 128k 1.0 1.0 max_tokens: 4096; frequency_penalty: 0.0;
Gemini Pro gemini-1.0-pro 32k 0.9 1.0 maxOutputTokens: 2048;
Llama 2 7B llama-2-7b-chat 4k 1.0 0.9 max_seq_len: 4096;
Llama 2 70B llama-2-70b-chat 4k 1.0 0.9 max_seq_len: 4096;
DeepSeek-R1 deepseek-r1-250528 128k 1.0 0.7 max_tokens: 4096; frequency_penalty:0.0;
@@ 2.3.1 LLM Selection
To fully evaluate the performance of LLMs in security defect detection, we firstly selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December 2023. Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through the API. For open-source LLMs, we deployed the smallest and largest chat versions of models in the Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs, each with 24 GB of VRAM. Given the rapid development of LLMs and the emergence of reasoning-optimized models, we additionally employed DeepSeek-R1 and accessed its updated version released on May 28, 2025 through the API. To realistically reflect on the practical capabilities of LLMs in security defect detection, we used the default hyperparameter settings throughout our experiments. Some studies mitigate the non-determinism of LLMs by setting the temperature to zero (atil2024llm; blackwell2024towards). In contrast, we refrained from doing so, as this may adversely affect the performance of the LLMs (song2024good) and may not guarantee determinism (ouyang2025empirical; astekin2024exploratory). We retained the default temperature settings and quantified the non-determinism of LLMs as one aspect of performance evaluation in RQ1.4, rather than artificially suppressing such non-determinism via hyperparameter tuning. The versions, context windows, and parameter settings of LLMs used in our study are summarized in Table 1.
Notably, GPT-3.5 and GPT-4 can be accessed through both API and the ChatGPT platform. During our preliminary exploration stage, we compared the performances of GPT-4 and GPT-3.5 accessed through these two ways. The two models provided by the ChatGPT platform performed better and were thus employed in our study. As the default parameters of the employed GPT-4 and GPT-3.5 were not publicly disclosed, the corresponding entries in Table 1 are indicated with ‘-’.